# The PLATO Mission

### v1

@@ 3.2.1 Brief summary of state of the art
Derived planetary mass-radius relationships and mean densities are important indicators of the nature of detected planets. To first order, the mean density-mass relationship can be used as an indicator for planet classification: [142] use it to reveal the turnover from small terrestrial and Neptune-like planets to gas giant planets and further for the transition to brown dwarfs. More frequently mass-radius diagrams are used to obtain a first indication on the nature of planets (starting with, e.g., [277, 3, 311], and many since). See [241] for a recent study of transiting planets up to 120 MEarth. Even though a unique one-to-one mapping to particular planetary internal structures and compositions is difficult just based on radius and mass data alone due to the inherent degeneracy of the problem (see, e.g., Valencia et al 298, Rogers and Seager 265, Dorn et al 101), these data do provide important constraints to formation models when ensemble properties become apparent.
@@
The situation on deriving internal structure and composition of planets improves when simplifying assumptions can be made like elemental abundances taken from host stars (e.g. Dorn et al 102, Brugger et al 52) or information from our knowledge of planets in the Solar System. Although this greatly reduces the various degeneracies, it is clear that many assumptions currently made are over-simplified. For example, the host star photospheric composition today may not reflect that of its planets [251] as additional processes including devolatilization trends [314] or diverse impact scenarios can alter the planet’s composition, whereas material falling into the star might affect the measured stellar composition. Therefore, planet characterisation models should not rely only on stellar constraints, in view of large error bars that prevent a definite conclusion. While degeneracies can likely never be completely resolved, they can be reduced when the primary parameters are precisely measured (as done with PLATO) and when combined with additional data, in particular providing information on the planetary atmospheres (e.g. with JWST, ARIEL, ELT, LUVOIR/HABEX (now the Habitable World Observatory, HWO), LIFE). A key factor will be to increase the available statistical sample of well-characterised small planets.
@@
A recent example of planet characterisation by combining density measurements with further atmospheric data are TRAPPIST-1 b and c. Emission photometry of planet b shows a bare rock surface with apparently no atmosphere [131], while planet c discards a thick CO2 atmosphere [320]. [2] use these measurements to constrain the Fe-to-Si ratio of TRAPPIST-1 b based only on the density of the planet, without making assumptions on the composition of the star, by performing a retrieval on the mass and radius using a two-layer interior structure model. PLATO will provide a wealth of data to further follow such approaches to planet characterisation.
@@
The case is particularly challenging when considering planets in between predominantly gaseous or rocky nature, the so-called mini-neptunes (or ´mini gas planets´, hence small gaseous planets) and earths to super-earths-like planets. If one assumes for simplicity that such planets would consist of four distinct layers (i.e. an iron core, a silicate layer, a water/ice layer and an atmosphere), several combinations of these layers of different inherent densities or mass fractions can result in the same mean planet density. Early studies of super-earths with little or no hydrogen atmosphere showed that planet radius measurements to better than 5 together with masses determined to better than 10, such as PLATO will provide, would allow us to distinguish between an icy or rocky composition (e.g. Valencia et al 298). Recent studies with more advanced interior models, allowing for additional phases and internal layers in the interior, confirmed these requirements on radius and mass precision when constraining the planets interior structure [241, 100, 30].
@@
Furthermore, depending on the internal energy and in turn on the age of a planet, large portions of the silicate layer may be molten, leading to water being dissolved in the magma ocean rather than a separation of the two layers, which would directly influence the planet’s observed density [100]. The internal energy further determines if the metallic core can efficiently separate from the mantle, which also influences the measured planet radius and hence density [106, 192]. The precise measurement of the age of the planets will help to assess the likelihood of different endmember scenarios, since cooling and differentiation are strongly time-dependent processes.
@@
Understanding the nature of exoplanets is a big puzzle to be resolved, requiring the combination of all available observational information. The role of PLATO in solving this puzzle is in the provision of exploring accurate planet parameters (including the ages) in a homogeneously processed sample of objects.
@@ 3.2.2 Expected planetary radius accuracy
PLATO requirements for planetary radius and mass are defined for a reference Earth-Sun scenario: an Earth-sized planet orbiting a G0V star as bright as mag at 1 AU (see Section 8). The required planet radius accuracy is 5% for this reference case. For the brightest targets (10 mag) a radius with 3% accuracy should be achieved. This corresponds to an accuracy for of 2%. We note that these requirements are in many ways the most difficult case and larger planets and/or smaller host stars will provide even better precisions, depending on respective signal-to-noise levels and stellar models.
For now, we assume that the required stellar radius accuracy of 2% is reached (see Section 4.4 for a discussion on PLATO’s capacity to deduce stellar parameters). Hence the precisions on planetary radii discussed here provide accurate radii in terms of our terminology. Planetary radii are then obtained by fitting the shape of observed transit events in the processed light curves. Residual noise sources (instrumental or stellar) therefore could affect the achievable precision. In Section 11 we show that PLATO light curves will be dominated by white noise for stars between and 12 mag. Hence, PLATO instrumental noise effects are minimal for the P1 sample and even for the bright end of the P5 sample. What remains is a potential impact by residual stellar variability in the light curves. In a first step we neglect such an impact. See the literature overview of detailed studies made at the end of this section, showing that such an assumption is justified for the bright samples.
@@ 3.3 Constraints on planetary atmospheres
Although the PLATO mission is not primarily designed to study exoplanetary atmospheres, its light curves can nevertheless provide relevant information on atmospheric properties. The amplitude and shape of white-light orbital phase curves, for example, can provide a first constraint on atmospheric meridional transport, mass and albedo. Basic colour information (“red” and “blue”) from the broadband filters on the PLATO fast cameras (with fast read-out cadence) can constrain the bulk atmospheric composition from estimating the Rayleigh spectral absorption feature. Also, PLATO’s prime data products, namely planetary radius, mass, and age, are crucial for retrieval and interpretation of atmospheric properties from spectroscopic data.
@@
Grenfell et al [132] found that planetary geometric albedos, as well as moderate-to-strong Rayleigh extinction can be detected for targets up to 25-100 pc distance based on simulated SNR for observations of nearby hot and ultra-hot Jupiters (UHJs) with the PLATO fast cameras. Their work suggests that initial constraints can be deduced from UHJ phase curve amplitudes for nearby targets closer than 25 pc, although this will be challenging. To illustrate this, Figure 8 shows a simulated phase curve for a hypothetical close-by (10 pc away) UHJ as would be observed by the PLATO fast cameras. From phase curve data on UJHs as shown in Figure 8 one can infer the dayside and nightside temperatures and thus the energy balance and intrinsic heat flow as well as planetary albedos, which are very important planet properties. Given that at present there are about only two dozen planets with observed phase curves, any one that PLATO can add will be valuable.
@@
Regarding warm sub-Neptunes and Super-Earths, the Grenfell et al [132] study suggested that basic atmospheric bulk compositions and haze properties in the upper atmosphere can be distinguished for some nearby (10 pc) favoured targets, although this will likely be a challenging task. Carrión-González et al [63, 64] point out that having prior information from PLATO on the planetary radius is an important input for model studies constraining atmospheric clouds and composition from direct imaging, since being able to fix the radius helps to disentangle degeneracies between planet radius, cloud layers and major atmospheric absorbers when studying planets in reflected light. [297, 274] suggest that PLATO measurements of planetary radii for hot Super-Earths with giant steam atmospheres will allow us to pinpoint and characterise runaway greenhouse regimes with strongly inflated atmospheres [see also 49]. [239] offer a way to distinguish lighter, more extended Super-Earth atmospheres (outgassed by more reducing mantles [192]) from more compact atmospheres outgassed by more oxidised mantles from the high-precision PLATO measurements of the planetary radius.
@@ 3.4 Constraints on planet formation and evolution
These processes include protoplanetary disc evolution [188], the growth of planetary embryos through pebbles and/or planetesimals [222, 165], the accretion of gas onto forming planets [e.g. 81, 185, 230], and disc-driven migration [178]. Although much progress has been made in the past decade on these topics, some key physical mechanisms and their relative importance remain poorly understood. For example, a variable mass budget available in pebbles could regulate the formation and migration of either Earth-like or super-Earth-like planets in the terrestrial zone [186]. Disc evolution in turn regulates migration rates and resonant-trapping of planet systems [169, 29, 153]. Finally, the composition of planets is a complex product of the full planetary growth process, starting from the growth and drift of primordial dust grains [e.g. 166, 275]. Models also exist of processes that act on longer timescales, such as tidal evolution, core-powered and photoevaporative gas envelope loss, and the role of stellar cluster membership in shaping planetary systems [e.g. 244, 319, 264, 189].
Progress will be made through comparison between these models and the PLATO-observed planet population which will then shed light on the relative importance of these processes, and will point to areas where model improvements are required, either through the inclusion of hitherto neglected physical effects or through improved understanding and modelling of fundamental planet formation processes [e.g. 84, 44, 108, 160, 253]. In addition to examining the global planet population, the models will explore important correlations in the data, for example between host star metallicity or stellar type and the emerging architectures of planetary systems, shedding additional light on the factors that control planet formation [196].
PLATO will not only find planets around single main-sequence stars, but is expected to find substantial numbers of more exotic planets, such circumbinary planets, that provide more extreme environments that can act as stringent tests of planet formation theories. For example, allowing discrimination between in situ scenarios and those that require migration [e.g. 250]. Development of population synthesis models of these types of systems [85] will allow comparison with PLATO data, constraining theories of planet formation in diverse environments.
@@ 4.1 State of the art of seismic stellar characterisation
Starting with the Sun, two decades of observations have demonstrated that solar-like seismic constraints are the most powerful tool to derive precise stellar masses, radii, densities, and ages, provided that high-quality seismic parameters are available [for reviews see 70, 74, 119, 279]. However, this requires short-cadence (less than 1 min for dwarfs and subgiants), ultra-high photometric precision (at the level of parts-per-million or ppm) and nearly-uninterrupted long-duration (from months to years) monitoring as best provided by space observations.
Beyond the precise characterisation of specific stars, seismology allows us to probe the physical interiors. In this way, it allows the identification of shortcomings in stellar modelling, to guide the theoretical developments for improvements in the physical description and then validate those improvements. This in turn leads to a more accurate characterisation of all stars – in particular their age-dating, not only for those with seismic features. Compared to classical modelling without asteroseismology, higher accuracy is achieved for the properties and parameters measured for a given star from the use of seismic constraints and more realistic stellar modelling [283, 236] or through the use of seismically calibrated relationships such as gyrochronology [12, 306, 137, and references therein] and abundance ratios – age relations [233, 223].
@@ 4.1.1 Ultra-precise photometric seismology: the space era
After the proof of concept space missions – either by accident [WIRE: 59] or dedicated (MOST: [313] – the French-led CoRoT satellite [2006 – 2009; 21, 18] really opened the space-seismic era with the detection of solar-like oscillations in about twelve bright dwarfs and subgiants [213] and thousands of red-giant stars [97, 214]. This has led to the first detailed asteroseismic studies of solar-like oscillations in main-sequence stars, as well as to the new field of red giant seismology [e.g. 20].
The NASA space mission Kepler [2009 – 2018; 47] monitored more than 150000 main sequence stars, of which about 2600 dwarfs and subgiants were observed for one month in short-cadence mode (58.85 sec). These measurements turned detailed seismic investigations into a reality for large samples of dwarfs [126, 71]. Overall, the nominal Kepler mission led to the detection of solar-like oscillations for more than 600 dwarfs and subgiants. Most of them were observed for only one month while about 150 stars were observed for longer, up to 4 years. The so-called Kepler Legacy sample of 66 dwarfs revealed the major power of asteroseismology of Sun-like stars [198, 283]. A few dozen of them host planets and these are now among the most precisely characterised host stars [156, 282, 43, 141].
The NASA TESS space mission launched in 2018 [259] is currently in operation. It provides lower-quality photometry than the Kepler mission but focuses on bright stars () and carries out a nearly-all-sky survey. The observing time is short, 27 days for most stars, but it reaches up to 352 days in the continuous viewing zone. During its 2 year nominal mission, TESS had a short-cadence mode of 2 min. During the first extended mission an additional 20 second mode was operational, which is better suited for seismic analyses as advocated by [159]. As demonstrated by the first results, TESS is capable of detecting solar-like oscillations with high amplitudes and therefore suitable to study seismology of subgiants and red giants [e.g., 61, 154, 73].
@@
4.1.2 Ultra-precise seismic stellar characterisation with Kepler

### v2

@@ 3.2.1 Brief summary of state of the art
Derived planetary mass-radius relationships and mean densities are important indicators of the nature of detected planets. To first order, the mean density-mass relationship can be used as an indicator for planet classification: [155] use it to reveal the turnover from small terrestrial and Neptune-like planets to gas giant planets and further for the transition to brown dwarfs. More frequently mass-radius diagrams are used to obtain a first indication on the nature of planets (starting with, e.g., [301, 3, 335], and many since). See [262] for a recent study of transiting planets up to 120 MEarth. Even though a unique one-to-one mapping to particular planetary internal structures and compositions is difficult based on radius and mass data alone due to the inherent degeneracy of the problem (see, e.g., Valencia et al 322, Rogers and Seager 288, Dorn et al 109), these data nevertheless provide important constraints on formation models when ensemble properties become apparent.
@@
The situation on deriving internal structure and composition of planets improves when simplifying assumptions can be made like elemental abundances taken from host stars (e.g. Dorn et al 110, Brugger et al 57) or information from our knowledge of planets in the Solar System. Although this greatly reduces the various degeneracies, it is clear that many assumptions currently made are over-simplified. For example, the host star photospheric composition today may not reflect that of its planets [275] as additional processes including devolatilization trends [338] or diverse impact scenarios can alter the planet’s composition, whereas material falling into the star might affect the measured stellar composition. Therefore, planet characterisation models should not rely only on stellar constraints, in view of large error bars that prevent a definite conclusion. While degeneracies can likely never be completely resolved, they can be reduced when the primary parameters are precisely measured (as done with PLATO) and when combined with additional data, in particular providing information on the planetary atmospheres (e.g. with JWST, ARIEL, ELT, LUVOIR/HABEX (now the Habitable World Observatory, HWO), LIFE). A key factor will be to increase the available statistical sample of well-characterised small planets.
@@
A recent example of planet characterisation by combining density measurements with further atmospheric data are TRAPPIST-1 b and c. Emission photometry of planet b suggests a bare rock surface with apparently no atmosphere [143], while planet c is unlikely to have a thick CO2 atmosphere [344]. [2] use these measurements to constrain the Fe-to-Si ratio of TRAPPIST-1 b based only on the density of the planet, without making assumptions on the composition of the star, by performing a retrieval on the mass and radius using a two-layer interior structure model. PLATO will provide a wealth of data to further follow such approaches to planet characterisation.
@@
The case is particularly challenging when considering planets in-between a predominantly gaseous or rocky nature, the so-called mini-neptunes (or ´mini gas planets´, hence small gaseous planets) and also when considering masses from earths to super-earths. If one assumes for simplicity that such planets would consist of four distinct layers (i.e. an iron core, a silicate layer, a water/ice layer and an atmosphere), several combinations of these layers of different inherent densities or mass fractions can result in the same mean planet density. Early studies of super-earths with little or no hydrogen atmosphere showed that planet radius measurements to better than 5 together with masses determined to better than 10, such as PLATO will provide, would allow us to distinguish between an icy or rocky composition (e.g. Valencia et al 322). Recent studies with more advanced interior models, allowing for additional phases and internal layers in the interior, confirmed these requirements on radius and mass precision when constraining the planets interior structure [262, 108, 30].
@@
Furthermore, depending on the internal energy and in turn on the age of a planet, large portions of the silicate layer may be molten, leading to water being dissolved in the magma ocean rather than a separation of the two layers, which would directly influence the planet’s observed density [108]. The internal energy further determines if the metallic core can efficiently separate from the mantle, which also influences the measured planet radius and hence density [115, 210]. The precise measurement of planetary age will help to assess the likelihood of different endmember scenarios, since cooling and differentiation are strongly time-dependent processes.
@@
Understanding the nature of exoplanets is a significant puzzle to be resolved, requiring the combination of all available observational information. The role of PLATO in solving this puzzle is in the provision of exploring accurate planet parameters (including the ages) in a homogeneously processed sample of objects.
@@ 3.2.2 Expected planetary radius accuracy
PLATO requirements for planetary radius and mass are defined for a reference Earth-Sun scenario: an Earth-sized planet orbiting a G0V star as bright as mag at 1 AU (see Section 8). The required planet radius accuracy is 5% for this reference case. For the brightest targets (10 mag) a radius with 3% accuracy should be achieved. This corresponds to an accuracy for Rplanet/Rstar of 2%. We note that these requirements are in many ways the most difficult case and larger planets and/or smaller host stars will provide even better precisions, depending on respective signal-to-noise levels and stellar models.
For now, we assume that the required stellar radius accuracy of 2% is reached (see Section 4.4 for a discussion on PLATO’s capacity to deduce stellar parameters). Hence the precisions on planetary radii discussed here provide accurate radii in terms of our terminology. Planetary radii are then obtained by fitting the shape of observed transit events in the processed light curves. Residual noise sources (instrumental or stellar) therefore could affect the achievable precision. In Section 11 we show that PLATO light curves will be dominated by white noise for stars between and 12 mag. Hence, PLATO instrumental noise effects are minimal for the P1 sample and even for the bright end of the P5 sample. What remains is a potential impact by residual stellar variability in the light curves. In a first step we neglect such an impact (see the literature overview of detailed studies made at the end of this section, showing that such an assumption is justified for the bright samples).
@@ 3.3 Constraints on planetary atmospheres
Although the PLATO mission is not primarily designed to study exoplanetary atmospheres, its light curves can nevertheless provide relevant information on atmospheric properties. The amplitude and shape of white-light orbital phase curves, for example, can provide a first constraint on atmospheric meridional transport, mass and albedo. High precision white light photometry has already been performed to measure geometric albedos e.g. for the Hot Jupiter, HD209458b by Brandeker et al [52] and by Deline et al [105] for the Ultra Hot Jupiter, WASP-189b. Basic colour information (“red” and “blue”) from the broadband filters on the PLATO fast cameras (with fast read-out cadence) can constrain the bulk atmospheric composition from estimating the Rayleigh spectral absorption feature. Also, PLATO’s prime data products, namely planetary radius, mass, and age, are crucial for retrieval and interpretation of atmospheric properties from spectroscopic data.
@@
Grenfell et al [144] found that planetary geometric albedos, as well as moderate-to-strong Rayleigh extinction can be detected for targets up to 25-100 pc distance based on simulated SNR for observations of nearby hot and ultra-hot Jupiters (UHJs) with the PLATO fast cameras. Their work suggests that initial constraints can be deduced from UHJ phase curve amplitudes for nearby targets closer than 25 pc, although this will be challenging. To illustrate this, Figure 8 shows a simulated phase curve for a hypothetical close-by (10 pc away) UHJ as would be observed by the PLATO fast cameras. From phase curve data on UJHs as shown in Figure 8 one can infer the dayside and nightside temperatures and thus the energy balance and intrinsic heat flow as well as planetary albedos, which are very important planet properties. Given that at present there are about only two dozen planets with observed phase curves, any additions that PLATO can make will be valuable.
@@
Regarding warm sub-Neptunes and Super-Earths, the Grenfell et al [144] study suggested that basic atmospheric bulk compositions and haze properties in the upper atmosphere can be distinguished for some nearby (10 pc) favoured targets, although this will likely be a challenging task. Carrión-González et al [69, 70] point out that having prior information from PLATO on the planetary radius is an important input for model studies constraining atmospheric clouds and composition from direct imaging, since being able to fix the radius helps to disentangle degeneracies between planet radius, cloud layers and major atmospheric absorbers when studying planets in reflected light. [321, 298] suggest that PLATO measurements of planetary radii for hot Super-Earths with giant steam atmospheres will allow us to pinpoint and characterise runaway greenhouse regimes with strongly inflated atmospheres, yielding an empirical test for the habitable zone hypothesis [see also 51]. [259] offer a way to distinguish lighter, more extended Super-Earth atmospheres (outgassed by more reducing mantles [210]) from more compact atmospheres outgassed by more oxidised mantles from the high-precision PLATO measurements of the planetary radius.
@@ 3.4 Constraints on planet formation and evolution
These processes include protoplanetary disc evolution [207], the growth of planetary embryos through pebbles and/or planetesimals [245, 176], the accretion of gas onto forming planets [e.g. 86, 201, 253], and disc-driven migration [191]. Although much progress has been made in the past decade on these topics, some key physical mechanisms and their relative importance remain poorly understood. For example, a variable mass budget available in pebbles could regulate the formation and migration of either Earth-like or super-Earth-like planets in the terrestrial zone [202]. Disc evolution in turn regulates migration rates and resonant-trapping of planet systems [180, 29, 167]. Finally, the composition of planets is a complex product of the full planetary growth process, starting from the growth and drift of primordial dust grains [e.g. 177, 299]. Models also exist of processes that act on longer timescales, such as tidal evolution, core-powered and photoevaporative gas envelope loss, and the role of stellar cluster membership in shaping planetary systems [e.g. 267, 343, 287, 208].
Progress will be made through comparison between these models and the PLATO-observed planet population which will then shed light on the relative importance of these processes, and will point to areas where model improvements are required, either through the inclusion of hitherto neglected physical effects or through improved understanding and modelling of fundamental planet formation processes [e.g. 89, 43, 117, 172, 277]. Through the diverse composition of PLATO’s stellar sample, predicted and newly emerging demographic trends can be studied. In particular, the occurrence rate of small planets around cool M stars will be determined and put into context with models constraining their formation [215, 237, 64, 297]. PLATO will further shed light on the existence and occurrence rate of terrestrial planets with significant water content (e.g., [188]), thereby yielding important constraints on water delivery on Earth and other terrestrial planets. Studies of architectural trends within multiplanet systems will be facilitated by PLATO’s large expected planet yield, shedding additional light on the factors that control planet formation (e.g., [136, 240, 241].)
PLATO will not only find planets around single main-sequence stars, but is expected to find substantial numbers of more exotic planets, such circumbinary planets, that provide more extreme environments that can act as stringent tests of planet formation theories. For example, allowing discrimination between in situ scenarios and those that require migration [e.g. 273]. Development of population synthesis models of these types of systems [90] will allow comparison with PLATO data, constraining theories of planet formation in diverse environments.
@@ 4.1 State of the art of seismic stellar characterisation
Starting with the Sun, two decades of observations have demonstrated that solar-like seismic constraints are the most powerful tool to derive precise stellar masses, radii, densities, and ages, provided that high-quality seismic parameters are available [for reviews see 76, 79, 130, 303]. However, this requires short-cadence (less than 1 min for dwarfs and subgiants), ultra-high photometric precision (at the level of parts-per-million or ppm) and nearly-uninterrupted long-duration (from months to years) monitoring as best provided by space observations.
Beyond the precise characterisation of specific stars, seismology allows us to probe the physical interiors. In this way, it allows the identification of shortcomings in stellar modelling, to guide the theoretical developments for improvements in the physical description and then validate those improvements. This in turn leads to a more accurate characterisation of all stars – in particular their age-dating, not only for those with seismic features. Compared to classical modelling without asteroseismology, higher accuracy is achieved for the properties and parameters measured for a given star from the use of seismic constraints and more realistic stellar modelling [306, 256] or through the use of seismically calibrated relationships such as gyrochronology [13, 330, 150, and references therein] and abundance ratios – age relations [255, 246].
@@ 4.1.1 Ultra-precise photometric seismology: the space era
After the proof of concept space missions – either by accident [WIRE: 65] or dedicated (MOST: [337] – the French-led CoRoT satellite [2006 – 2009; 22, 19] really opened the space-seismic era with the detection of solar-like oscillations in about twelve bright dwarfs and subgiants [234] and thousands of red-giant stars [102, 235]. This has led to the first detailed asteroseismic studies of solar-like oscillations in main-sequence stars, as well as to the new field of red giant seismology [e.g. 21].
The NASA space mission Kepler [2009 – 2018; 49] monitored more than 150000 main sequence stars, of which about 2600 dwarfs and subgiants were observed for one month in short-cadence mode (58.85 sec). These measurements turned detailed seismic investigations into a reality for large samples of dwarfs [137, 77]. Overall, the nominal Kepler mission led to the detection of solar-like oscillations for more than 600 dwarfs and subgiants. Most of them were observed for only one month while about 150 stars were observed for longer, up to 4 years. The so-called Kepler Legacy sample of 66 dwarfs revealed the major power of asteroseismology of Sun-like stars [218, 306]. A few dozen of them host planets and these are now among the most precisely characterised host stars [169, 305, 42, 154].
The NASA TESS space mission launched in 2018 [282] is currently in operation. It provides lower-quality photometry than the Kepler mission but focuses on bright stars () and carries out a nearly-all-sky survey. The observing time is short, 27 days for most stars, but it reaches up to 352 days in the continuous viewing zone. During its 2 year nominal mission, TESS had a short-cadence mode of 2 min. During the first extended mission an additional 20 second mode was operational, which is better suited for seismic analyses as advocated by [171]. As demonstrated by the first results, TESS is capable of detecting solar-like oscillations with high amplitudes and therefore suitable to study seismology of subgiants and red giants [e.g., 67, 168, 78].
@@
4.1.2 Ultra-precise seismic stellar characterisation